OpenStudy (anonymous):

Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1620. Part A: Create an equation that will determine the number of days x on the trip. Part B: Solve this equation justifying each step with an algebraic property of equality. Part C: How many days did Donna spend on this trip?

OpenStudy (anonymous):

Well, she drove 200 miles, at a pay of $0.60 per mile travelled. 200 * 0.6 = 120 She got payed $120 for her 200 miles traveled. Since she gets payed $50 per day, and she got paid $1500 for her food and lodging, she traveled for 30 days as 1500/50 = 30 (Part C answer) If you were to determine her total pay: Using variables x for: number of days spent on trip Y: number of miles traveled. Total pay = 50x + 0.6y If you'd rather have only the equation for the number of days spent on her trip: X = (1620- 0.6(200))/50. (Part A answer) Solving for x: X = (1620 - 0.6(200))/50 X = (1620 - 120)/50 X = (1500)/50 X = 30 (Part B answer)

OpenStudy (anonymous):

Well damn thank you :)

OpenStudy (anonymous):

Sure thing :)